# Mechanism of magnetic phase transition in correlated magnetic metal: insight into itinerant ferromagnet Fe3-δGeTe2

### v1

@@ Results
Another critical issue is the emergence of heavy fermion behavior within ferromagnetic order [22], as the two often contradict each other. DFT+DMFT calculations in Fig. 3e show similar peaks near the Fermi level compared to STM measurements [22, 23]. As shown in Fig. 3d and e, DFT+DMFT calculations illustrate some peaks just above the Fermi level, indicated by a red circle. As experiments suggest, these peaks possibly originate from the Kondo resonance [22, 23]. According to our calculations, these peaks come from the hole-like flat quasiparticle bands located at the $\Gamma$ point, promoted by the rapid decrease of spin fluctuations in ferromagnetic order. Furthermore, as temperature decreases, the quasiparticle bands at the $\Gamma$ point show strong hybridization with other relatively itinerant bands, where the partial DOS of Fe1, Fe2, and Ge atoms have similar shapes and peaks positions over a wide range (Supplementary Information Fig.3). These $\Gamma$-centered hole-like bands have a small Fermi velocity in Fig. 3c. Based on this evidence, we conclude that Hund’s coupling can promote the formation of ferromagnetic order and heavy fermion behavior simultaneously. This mechanism demonstrates, for the first time, the complex competition and cooperation of magnetic order, Hundness, Mottness, and heavy fermion behavior in a detailed manner.
@@ Discussion
In strongly correlated physics, various exotic phenomena are closely related to the local moments of correlated atoms. One such outstanding material is the strongly correlated ferromagnetic metal FGT. However, the origin of magnetism in itinerant ferromagnets has been a controversial topic for several decades. As shown in Fig. 4, the Heisenberg model is used to describe the origin of magnetism in insulators [15]. In contrast, the Stoner model, based on the mean-field approximation, provides a phenomenological description of spin splitting in magnetic metals [16]. However, the Stoner model has been proven inadequate for these correlated systems [27, 30]. Even though Moriya and Kawabata attempted to consider spin fluctuations and correlation effects through self-consistent renormalization (SCR) theory, a detailed physical picture still needs to be provided [45, 20]. Our magnetic DFT+DMFT calculations aim to resolve this problem from the many-body dynamical correlation point of view. In ferromagnetic metals, Hund’s coupling can produce collective excitation modes at low temperatures and create renormalized flat quasiparticle bands through dynamical correlation processes. Meanwhile, the spectral weight transfer occurs during magnetic phase transition. Additionally, the formation of ferromagnetic long-range order, in turn, promotes quasiparticle formation through positive feedback by rapidly reducing spin fluctuations in the microscopic mechanism. Ultimately, this leads to the simultaneous occurrence of spectral weight transfer, long-range magnetic order, and even other competing orders, such as heavy fermion behavior.
@@
In summary, through systematic magnetic DFT+DMFT calculations on FGT, we have uncovered a new microscopic description of electronic structures in correlated ferromagnetic metals based on dynamical correlation effects. The Hund’s coupling results in long-range ferromagnetic order accompanied by spectral weight transfer, rather than the spin splitting described by the Stoner model. This long-range magnetic order suppress spin fluctuation rapidly, leading to the formation of renormalized quasiparticle bands. Our work also show that heavy fermion behavior can be promoted by long-range magnetic order in itinerant magnets, through flat quasiparticle bands around the Fermi level further hybridizing with other itinerant bands at lower temperatures. In addition, the two Fe sites exhibit different physics, with Fe1 characterized by Mottness and Fe2 by Hundness, induced by the diversity of their atomic environment. Overall, our research indicates that multi-site itinerant magnets present a new frontier for discovering new quantum states and their competitions.
@@ Methods
The DFT+DMFT calculations were employed to treat the electronic correlations of Fe-$d$ orbitals. Full-potential linearized augmented plane-wave method implemented in the WIEN2k package was used in DFT+DMFT [50]. We employed the hybridization expansion continuous-time quantum Monte Carlo method as the impurity solver [51, 52] and applied the exact double-counting scheme [42]. The correlation parameters were set $U=$ 5 eV and $J=$ 0.8 eV, as suggested from earlier work [24]. We used the exact double-counting scheme, and it can get the proper occupied number of Fe atoms, consistent with experiments [17, 19]. To account for the effect of Fe vacancy, we modified our DFT+DMFT calculations by changing the number of electrons. Previous research has confirmed that changing electron numbers leads to similar results in constructing supercells with Fe2 vacancies [44, 53]. In order to perform magnetic DFT+DMFT calculations, we consider 10 Fe-$d$ orbitals and break the symmetry with a change in the real-part energy of the initial self-energy of both two types of Fe atoms.
@@ Acknowledgements
The authors thank Yingying Cao, Beilei Liu, Huican Mao and Kun Zhai for the fruitful discussions. This work is supported by the National Natural Science Foundation of China (Grant Nos. U2230401, 12204033, 52371174), the National Key Research and Development Program of China (Grant No. 2021YFB3501503), the Foundation of LCP, the Fundamental Research Funds for the Central Universities (Grant No. FRF-TP-22-097A1) and the Young Elite Scientist Sponsorship Program by BAST (Grant No. BYESS2023301). We thank the Tianhe platforms at the National Supercomputer Center in Tianjin.
@@
F. T. and H. S. conceived the idea and supervised the project; Y. X. and Y. W. performed the calculations; X. J., Y. L. and H. L. participated in the discussions. Y. X. wrote the paper with input from all authors. All authors contributed to the interpretation of the data and revised the paper. Y. X. and Y. W. contributed equally to this work.

### v2

@@ Results
Another critical issue is the emergence of heavy fermion behavior within ferromagnetic order [22], as the two often contradict each other. Measurements obtained from scanning tunneling microscopy (STM) reveal the formation of two peaks at -35 meV and -170 meV below the Fermi level, coinciding with the transition to ferromagnetic order. Additionally, a shoulder-like peak, identified as the Kondo peak, is observed at approximately 20 meV above the Fermi level at lower temperatures [22, 23]. In our DFT+DMFT calculations for Fe2.82GeTe2, the presence of low-temperature ferromagnetic order significantly reduces spin fluctuations, resulting in sharp quasiparticle flat bands and corresponding prominent peaks in the DOS, as illustrated in Fig. 3d. Notably, a distinct peak appears at around 20 meV above the Fermi level, indicated by the red circle in Figs. 3d and 3e. The emergence of these distinct peaks in the DOS near the Fermi level under ferromagnetic order is consistent with experimental observations.
Although the peak observed above the Fermi level aligns with the position found in experiments, its height is not sufficiently pronounced. Our calculations indicate that these peaks originate from the hole-like flat quasiparticle bands located at the $\Gamma$ point. Furthermore, as temperature decreases, the quasiparticle bands at the $\Gamma$ point show strong hybridization with other relatively itinerant bands, where the partial DOS of Fe1, Fe2, and Te atoms have similar shapes and peaks positions over a wide range (See Supplementary Note 7). Moreover, at low temperatures, the imaginary parts of the real-frequency hybridization functions exhibit a substantial increase near the Fermi level, particularly for the spin-down channel of the Fe1-$\rm e_{g}^{\prime}$ orbital (See Supplementary Note 8). Besides, these $\Gamma$-centered hole-like bands have a small Fermi velocity in Fig. 3c. Based on this evidence, we conclude that the possible hybridization between these well-defined flat bands and other itinerant bands may be the mechanism responsible for the heavy fermion behavior observed at low temperatures, similar to that seen in $f$-electron heavy fermion materials. Additionally, the crossover temperature for heavy fermion behavior is reported to be between 110 and 150 K based on experimental findings [22]. The large specific heat coefficient has been measured at lower temperatures [20]. And the Kondo peak can be distinctly identified in STM experiments at 6 K [22]. However, in our DFT+DMFT calculations, we have to consider ferromagnetism and the presence of two different Fe atoms. The computations conducted at lower temperatures will exceed our computational capabilities.
In contrast to $f$-electron heavy fermion systems, various mechanisms have been investigated in $d$-electron systems, such as geometrical frustration in LiV2O4 [49], symmetry-enforcement in CaCu3Ir4O12 [50]. In our study of FGT, Hund’s coupling promotes the formation of ferromagnetic order, which in turn reduces spin fluctuations and enhances the flat bands near the Fermi level. At lower temperatures, these well-defined flat bands may subsequently hybridize with other itinerant bands, leading to heavy fermion behavior. This proposed mechanism may offer a possible explanation of how heavy fermion behavior can be promoted by ferromagnetic order.
@@ Discussion
In strongly correlated physics, various exotic phenomena are closely related to the local moments of correlated atoms. One such outstanding material is the strongly correlated ferromagnetic metal FGT. However, the origin of magnetism in itinerant ferromagnets has been a controversial topic for several decades. As shown in Fig. 4, the Heisenberg model is used to describe the origin of magnetism in insulators [15]. In contrast, the Stoner model, based on the mean-field approximation, provides a phenomenological description of spin splitting in magnetic metals [16]. However, the Stoner model has been proven inadequate for these correlated systems [27, 30]. Our magnetic DFT+DMFT calculations aim to resolve this problem from the many-body dynamical correlation point of view. In ferromagnetic metals, the dual nature of $d$-electrons, arising from correlation effects, plays a crucial role. As Hund’s coupling can promote long-range ferromagnetic order, the opposing tendency in spectral weight transfer between the Hubbard bands in the two spin channels fully demonstrates the localized nature of $d$-electrons. Concurrently, the establishment of long-range ferromagnetic order enhances the formation of quasiparticle flat bands near the Fermi level through a positive feedback mechanism, which rapidly reduces spin fluctuations, thereby reflecting the itinerant characteristics of $d$-electrons. Ultimately, this leads to the simultaneous occurrence of spectral weight transfer, long-range magnetic order, and even other competing orders, such as heavy fermion behavior.
As FGT is a layered two-dimensional vdW material, non-local correlation effects may also play a significant role. Particularly in the vicinity of the magnetic transition, long-wavelength spin waves should not be overlooked [42]. A prior study stimulating the Curie temperature in Fe and Ni demonstrates that the single-site nature of the DMFT approach fails to account for the reduction of $T_{c}$ [42]. This overestimation of $T_{c}$ is also evident in our calculations. Moreover, the inadequate mass renormalization observed in our results may also highlight the necessity of incorporating non-local fluctuations, as suggested by previous investigations of iron [51]. Therefore, the non-local correlation effect in FGT warrant further investigation.
@@
In summary, through systematic magnetic DFT+DMFT calculations on FGT, we have uncovered a systematic microscopic description of electronic structures in correlated ferromagnetic metals based on dynamical correlation effects. The Hund’s coupling results in long-range ferromagnetic order accompanied by spectral weight transfer between the two spin channels. This long-range magnetic order suppress spin fluctuation rapidly, leading to the formation of renormalized quasiparticle bands. Our work also show that heavy fermion behavior can be promoted by long-range magnetic order in itinerant magnets, through flat quasiparticle bands around the Fermi level further hybridizing with other itinerant bands at lower temperatures. In addition, the two Fe sites exhibit different physics, with Fe1 characterized by Mott physics and Fe2 by Hund physics, induced by the diversity of their atomic environment. Overall, our research indicates that multi-site itinerant magnets present a new frontier for discovering new quantum states and their competitions.
@@ Methods
The DFT+DMFT calculations were employed to treat the electronic correlations of Fe-$d$ orbitals. Full-potential linearized augmented plane-wave method implemented in the WIEN2k package was used in DFT+DMFT [56]. We employed the hybridization expansion continuous-time quantum Monte Carlo method as the impurity solver [57, 58] and applied the exact double-counting scheme [46]. The correlation parameters were set $U=$ 5 eV and $J=$ 0.8 eV, as suggested from earlier work [24]. We used the exact double-counting scheme, and it can get the proper occupied number of Fe atoms, consistent with experiments [17, 19]. We have examined other double-counting schemes, including the nominal and fully localized limit (FLL) schemes. Although the occupation numbers of the Fe atoms vary with different double-counting schemes, the primary finding of spectral weight transfer in our study remains consistent. The occupation numbers of different schemes are presented in Supplementary Note 5. To account for the effects of Fe vacancy and reduce computational costs, we employed hole doping simulations by decreasing the total number of electrons in accordance with the concentration of Fe vacancies, as Fe deficiency induces hole doping in this material. Previous research has confirmed that changing electron numbers leads to similar results in constructing supercells with Fe2 vacancies [48, 59]. In order to perform magnetic DFT+DMFT calculations, we consider 10 Fe-$d$ orbitals and break the symmetry with a change in the real-part energy of the initial self-energy of both two types of Fe atoms.
@@ Acknowledgements
The authors thank Yingying Cao, Beilei Liu and Kun Zhai for the fruitful discussions. This work is supported by the National Natural Science Foundation of China (Grant Nos. U2230401, 12204033, 52371174), the National Key Research and Development Program of China (Grant No. 2021YFB3501503), the Foundation of LCP, the Fundamental Research Funds for the Central Universities (Grant No. FRF-TP-22-097A1), the State Key Lab of Advanced Metals and Materials (Grant No. 2022Z-13) and the Young Elite Scientist Sponsorship Program by BAST (Grant No. BYESS2023301). We thank the Tianhe platforms at the National Supercomputer Center in Tianjin. Numerical computations were also performed on Hefei advanced computing center.
@@
F.Y.T. and H.F.S. conceived the idea and supervised the project; Y.J.X. and Y.C.W. performed the calculations; X.T.J., Y.L. and H.F.L. participated in the discussions. Y.J.X.,F.Y.T. and H.F.S. wrote the paper with input from all authors. All authors contributed to the interpretation of the data and revised the paper. Y.J.X. and Y.C.W. contributed equally to this work.
Competing interests
The authors declare no competing interests.
Data availability
Data are available upon request.